# Biological computation through recurrence

### 07 Feb 2024 v1

@@ Abstract
One of the defining features of living systems is their adaptability to changing environmental conditions. This requires organisms to extract temporal and spatial features of their environment, and use that information to compute the appropriate response. In the last two decades, a growing body or work, mainly coming from the machine learning and computational neuroscience fields, has shown that such complex information processing can be performed by recurrent networks. In those networks, temporal computations emerge from the interaction between incoming stimuli and the internal dynamic state of the network. In this article we review our current understanding of how recurrent networks can be used by biological systems, from cells to brains, for complex information processing. Rather than focusing on sophisticated, artificial recurrent architectures such as long short-term memory (LSTM) networks, here we concentrate on simpler network structures and learning algorithms that can be expected to have been found by evolution. We also review studies showing evidence of naturally occurring recurrent networks in living organisms. Lastly, we discuss some relevant evolutionary aspects concerning the emergence of this natural computation paradigm.
@@ I Introduction
The survival of any living system depends on its capacity to sense multiple signals in its surroundings, integrate those external signals, and activate an adequate response. Organisms need to derive information from both the spatial and temporal structure of these signals [1]. Temporal information, in particular, is crucial to anticipate changes in the environment. In these cases, having some memory mechanism is clearly beneficial. In the mammalian brain, for instance, processing the temporal structure of visual stimuli gives information on the direction and speed of objects, as well as on the duration of and time between events [2]. In the somatosensory system, the contribution of temporal encoding is not limited to dynamic aspects of the stimulus, such as motion detection, but is also involved in stationary aspects such as object and texture discrimination [3]. It has been postulated that recurrence due to extensive feedback can underlie the temporal information-processing abilities of cortical neural networks [4, 2].
The importance of processing temporal information is not limited to neural systems. Individual, non-neural cells such as bacteria are able to relate groups of events that tend to occur together or in a specific order, as a form of associative memory [5]. For example, Escherichia coli prepares itself for low oxygen conditions when it detects a rise in temperature, an indication that it has been ingested by a mammal [6]. Similarly, Candida albicans cells, a fungal pathogen, upregulates oxidative stress resistance genes upon detection of an increase in glucose levels (which indicates they have entered a host), preemptively reacting to its immune system [7].
Apart from this associative memory capacity, recent studies indicate that microorganisms display both short- and long-term memory. For instance, the stress response of Bacillus subtilis relies not solely on its present growth conditions but also on its past [8]. The way in which this record of previous history —i.e. memory— is integrated and stored in bacteria cells is the subject of much interest. It has been postulated that this adaptability relies in general on a complex network of interactions between genes and proteins, frequently at the level of transcriptional regulation [9, 10]. These types of networks are called gene regulatory networks.
@@
The key characteristic of RC is that it separates memory encoding and prediction in different network substructures, as depicted in Fig. 1.
Figure 1: Structural and functional organization of the gene regulatory network of Escherichia coli. The genes represented in green form a recurrent subgraph with cyclic paths (the reservoir) that can maintain a record of the recent history in its dynamics. The genes in red constitute a directed acyclic subgraph (the readout) that interprets the information encoded in the reservoir state. The network architecture shown here was obtained from the Ecocyc database [15]. The representation is adapted from Gabalda-Sagarra, Carey, and Garcia-Ojalvo [11].
@@
    It acts as a nonlinear high-dimensional expansion of the input signal(s).
@@
    It allows the network to retain information for a certain time, giving memory to the system.
The second substructure, the readout, is a feedforward architecture located downstream of the reservoir (see red nodes in Fig. 1). The readout layer uses the history record encoded in the state of the reservoir, combined with the current input, to make a prediction or classification. This separation of roles allows to focus the training process (i.e. the modification of the network’s synaptic strengths so that the desired output is produced) solely on the readout [16].
In the classical RC approach, also known as an echo-state network (ESN), the reservoir is generated randomly, and only the readout is trained. Since feedforward architectures are much easier to train than recurrent ones, the RC framework has the computational power of a recurrent network and the ease of training of a feedforward one [2]. RC has demonstrated outstanding performance across various benchmark tasks [17, 18, 19].
@@ II.1 Basic description
It is worth noting that for the ESN approach to work, the network needs to be appropriately tuned. Specifically, the recurrent reservoir should satisfy the so-called echo state property: its state $\mathbf{x}_{t}$ should be uniquely defined by the fading memory of the input $\mathbf{u}_{t}$ [17]. In other words, for a long enough input $\mathbf{u}_{t}$, the reservoir state $\mathbf{x}_{t}$ should not depend on the initial conditions that were present before the input. It is empirically observed that the echo state property is obtained for any input if the spectral radius (i.e. the maximum absolute eigenvalue of the reservoir connectivity matrix $W$) is smaller than 1. Large spectral radius values can lead to reservoirs exhibiting multiple fixed points or periodic, even chaotic (when sufficient nonlinearity in the reservoir is reached) attractors, which violate the echo state property and prevent efficient training.
@@ II.2 Alternative training algorithms
To be operational, the learning algorithm must induce in the network a transition from chaotic to non-chaotic dynamics during training. This is precisely what the FORCE procedure achieves. Specifically, Sussillo and Abbott [23] show that the chaotic dynamics during training can be quenched by using either external feedback loops (Fig. 3a) or internal synaptic modifications (Figs. 3b and 3c).
Figure 3: Network architectures for the FORCE algorithm [23]. In all three cases, a reservoir (large circle with green nodes) drives a linear readout unit through weights that are modified during training (red arrows coming into the red node). Only connections shown in red are subject to modification. (a) Feedback to the reservoir is provided by the readout unit. (b) Feedback to the reservoir is provided by a separate feedback network (smaller circled network). The neurons of the feedback network are recurrently connected and receive input from the reservoir through synapses that are modified during training (in red). (c) Network with no external feedback. Instead, feedback is generated within the reservoir and modified by applying FORCE learning to the synapses internal to the network (red).
All three schemes lead to the generation of intricate yet controlled outputs. Driven by any of these two strategies, the key for suppressing chaos is making strong and rapid synaptic modifications. In this regard, FORCE learning works in a manner distinct from traditional neural network training. Typically, conventional training involves a series of adjustments aimed at gradually diminishing initially large errors in the network output. In contrast, the FORCE algorithm attains a large reduction in the error’s size at the time of the first weight update, and then maintains small errors throughout the training process. In order to achieve this, the method makes use of the recursive least-squares (RLS) algorithm [24] as the update rule for the trained weights (see Sussillo and Abbott [23] for details). Finally, by the end of the training period, modification is no longer needed, and the network can generate the desired output autonomously.
Lastly, it is worth emphasizing that the FORCE algorithm allows us to train not only the readout weights of the network, such as in the original RC framework (Fig. 3a), but also RNNs with the architectures shown in Figs. 3b and 3c, in which modifications are not restricted to the readout weights, but instead include internal synaptic modifications. These architectures match biological circuits better than that of Fig. 3a, where feedback comes directly (and exclusively) from the readout unit. By enabling training of different architectures, the FORCE algorithm extends the original RC framework, and gives more flexibility and biological feasibility to the concept of computation through recurrence.
@@ II.3 Using dynamical systems theory to understand recurrence-based computations
Since supervised training algorithms specify the computation to perform without constraining the network on how to implement this computation, the precise manner in which trained RNNs realize their target functions remains an open question. In essence, the resulting networks have predominantly been regarded as black boxes. This differs from network models deliberately designed to implement specific established mechanisms, such as those proposed, for instance, by Hopfield [25] and Wang [26]. One way of “opening the black box”, proposed by Sussillo and Barak [27], is to treat the network as a nonlinear dynamical system.
Sussillo and Barak [27] studied not only the existence of fixed points (both stable and unstable) and the linearized dynamics around them, but also the linearization in areas of phase space that are not true fixed points, but merely points in phase space through the system flows very slowly (slow points). They also present a simple optimization technique for locating such regions. Linearization around each one of the fixed and slow points provides a mechanistic description of how high-dimensional RNNs implement their computations [27]. This last point is key in order to suggest hypotheses for how the task might be solved by the biological network. Therein lies the importance of this dynamical-system based approach.
Identifying the fixed and slow points of an RNN requires optimization tools that rely on hard-coded analytic derivatives, e.g. of the hyperbolic tangent function (see e.g. Section 3.2 of Sussillo and Barak [27] for details). Algorithms to find fixed points have been developed for simple (Elman) RNNs [28], and Hopfield-like RNN architectures [27, 29], but can be inconvenient for complicated models, or for studies that involve a combination of different RNN models. In this sense, Golub and Sussillo [30] introduced an open-source Tensorflow toolbox for finding fixed points and linearized dynamics in arbitrary RNN architectures. This software profits from TensorFlow’s highly efficient automatic differentiation back-end to execute the necessary optimizations and compute Jacobians, essential for the identification and characterization of fixed points.
@@ III.1 The mammalian cerebral cortex
Cortical circuits were inspirational to Maass, Natschläger, and Markram [14] when conceiving LSMs. It has been proposed that cortical networks are likely to operate following an RC paradigm, or at least that they incorporate most RC principles. One indication of this is the fact that neurons vary morphologically and functionally throughout the cortex. Heterogeneity in neural characteristics has been observed in cortical neurons, the retina, and the primary visual cortex, as well as in networks cultured in vitro from dissociated neurons [31, 32, 33, 34, 35, 36]. This diversity of neural components results in very heterogeneous dynamics, which can be exploited by RC-like computation. In addition, cortical networks present a highly recursive nature [37].
Nonetheless, discerning the computational basis of cortical circuits is still today an open question in neuroscience, and efforts to clarify to what end recurrence is exploited in cortical circuits are under way [31, 32]. In what follows, we review related studies on cortical regions, including prefrontal and visual areas.
@@
An early RC-type model of the cortico-striatal system was proposed by Dominey, Arbib, and Joseph [38] to understand the mechanism of context-dependent oculomotor (eye movement) behavior [39]. The model consists of two sub-networks, based on the known cortical/subcortical anatomy of the primate oculomotor system. The first sub-network corresponds to the prefrontal cortex, which is associated with the planning of cognitive behavior, personality expression and decision making, besides moderating social behavior [40]. The second sub-network corresponds to the striatum, which is a subcortical area.
In the model of Dominey, Arbib, and Joseph [38], the prefrontal sub-network has fixed recurrent connections between its neurons (reservoir); while connections from the prefrontal cortex to the striatum are modifiable (readout) [41]. The neural activity in the prefrontal cortex network generates unique patterns of activity in response to the sequential visual inputs. The striatum associates these state-encoding patterns with the correct outputs, which represent the oculomotor movements.
Later, Dominey and Ramus [42] extended this model to the context of language processing in infants. Specifically, the authors showed that the previously mentioned cortico-striatal model (called a temporal recurrent network) simulates the capabilities of babies to learn the sequential and rhythmic structure of language. Indeed, the recurrent neural network is capable of representing serial and temporal structure with no pre-exposure, and through exposure these internal representations can become bound to behavioural responses. A modified version of the model, with an additional recurrent network with short-term memory, was however needed to reproduce the performance of babies in extracting abstract structure (or a hidden rule) from sequential inputs. These works illustrate how inherent representational abilities of the serial and temporal structure of language might emerge from a shared neural architecture, separate from what is needed for representing abstract structures.
The cortico-striatal model was further combined with neurophysiological models of language processing to learn grammatical constructions in sentence processing [43, 44] and with an RC framework to improve the performance of the learning algorithm [45]. The detailed history and the developments of the cortico-striatal model was reviewed by Dominey [46].
@@
The visual cortex is the part of the cerebral cortex that processes visual information. It is divided into the primary visual cortex (early visual areas) and visual association cortex (higher-order visual areas) [47]. Nikolić et al. [33] obtained experimental evidence showing that early visual areas in cats have fading memory properties, as demanded by LSMs. Specifically, in vivo spiking activities of 60–100 neurons were obtained for a sequence of different visual stimuli (letters). The authors showed that the neuronal responses contained information not only about the letter being presented but also about the preceding letters. In other words, the population activity encodes both the current and previous stimuli. In fact, the identity of the different visual stimuli (in this case letters) could be recovered with high reliability from the simultaneously recorded spike trains of neurons (corresponding to reservoir states) by a simple linear classifier emulating linear integrate-and-fire readout neurons [33].
The new perspective that the primary visual cortex can perform time-dependent computation for sequential inputs based on memory mechanisms contrasts with the conventional one that the brain executes frame-by-frame computation based on memory-less hierarchically organized feedforward architectures [48]. The limitation of the conventional viewpoint is that frame-by-frame computation cannot process dynamical visual scenes.
@@ III.2 Gene regulatory networks as reservoirs
Biological computation through recurrence does not necessarily need to arise in neural substrates, but can be supported by other biological networks. In this section, we review two example studies involving gene regulatory networks (GRNs). The first study looks into the ability of Escherichia coli to discriminate between complex temporal patterns of chemical inputs. The second one studies the memory capacities of the GRNs of five evolutionary distant organisms.
@@
Jones et al. [49] proposed that the gene regulatory network (GRN) of the bacterium Escherichia coli can act as a computational reservoir, receiving input signals from a sequence of chemical stimuli and acting upon a set of target proteins as an output. Specifically, the authors exposed a liquid culture of E. coli cells to sequences of inputs that represent different combinations of temperature and chemical conditions of the environment. Subsequently they studied the transcriptional state of the GRN using microarrays, to test whether the network was capable of making complex perceptual classification of the different inputs into desired classes. Numerical simulations corresponding to the above experimental design confirmed that the proposed RC framework was partially successful in an XOR problem, and showed good performance in a separation problem.
This work motivated an experimental study in which Didovyk et al. [50] genetically engineered bacterial cell cultures with the goal of realizing a variety of simple classifiers separating different chemical inputs. The outputs were then aggregated to form a more complex classifier.
@@
In the previous example, the GRN used as putative reservoir does not include any recurrences other than self-regulations of some sparsely interconnected nodes, and the system was tested with a task that does not require memory (classification task). Fully recurrent topologies are necessary, however, for processing time-dependent information. In this sense, Gabalda-Sagarra, Carey, and Garcia-Ojalvo [11] studied the computational capabilities of the GRNs of five evolutionary distant organisms (from bacteria to humans) in standard memory-demanding tasks. In all cases, they recognized a cyclic recurrent structure, constituted by a small fraction of the overall number of genes, which is crucial for dynamic encoding and information integration (reservoir). Meanwhile, the rest of the network forms a readout layer devoted to decode and interpret the high-dimensional dynamical state of the recurrent core. In that way, GRNs act as echo-state networks, providing the organism with a memory mechanism necessary to process temporal information.

### 25 Feb 2024 v2

@@ Abstract
One of the defining features of living systems is their adaptability to changing environmental conditions. This requires organisms to extract temporal and spatial features of their environment, and use that information to compute the appropriate response. In the last two decades, a growing body or work, mainly coming from the machine learning and computational neuroscience fields, has shown that such complex information processing can be performed by recurrent networks. Temporal computations arise in these networks through the interplay between the external stimuli and the network’s internal state. In this article we review our current understanding of how recurrent networks can be used by biological systems, from cells to brains, for complex information processing. Rather than focusing on sophisticated, artificial recurrent architectures such as long short-term memory (LSTM) networks, here we concentrate on simpler network structures and learning algorithms that can be expected to have been found by evolution. We also review studies showing evidence of naturally occurring recurrent networks in living organisms. Lastly, we discuss some relevant evolutionary aspects concerning the emergence of this natural computation paradigm.
@@ I Introduction
The survival of any living system depends on its capacity to sense multiple signals in its surroundings, integrate those external signals, and activate an adequate response. Not only the spatial, but also the temporal structure of these signals provide key information to the organisms [1]. Temporal information, in particular, is crucial to anticipate changes in the environment. In these cases, having some memory mechanism is clearly beneficial. In the mammalian brain, for instance, processing the temporal structure of visual stimuli gives information on the direction and speed of objects, as well as on the duration of and time between events [2]. In the somatosensory system, the contribution of temporal encoding is not limited to dynamic aspects of the stimulus, such as motion detection, but is also involved in stationary aspects such as object and texture discrimination [3]. It has been postulated that recurrence due to extensive feedback can underlie the temporal information-processing abilities of cortical neural networks [4, 2].
The importance of processing temporal information is not limited to neural systems. Non-neural individual cells such as bacteria possess the capability to associate groups of events that commonly co-occur or follow a particular sequence, thus exhibiting a form of associative memory [5]. As an illustration, Escherichia coli anticipates itself for low oxygen environments upon detecting an increase in temperature, which would signal ingestion by a mammal [6]. Similarly, Candida albicans cells, a fungal pathogen, upregulates oxidative stress resistance genes upon detection of an increase in glucose levels (which indicates they have entered a host), preemptively reacting to its immune system [7].
Apart from this associative memory capacity, recent studies indicate that microorganisms display both short- and long-term memory. For instance, the stress response of Bacillus subtilis relies not solely on its present growth conditions but also on its past [8]. The integration and storage of previous history, i.e. memory, in bacterial cells is a subject of considerable interest. It has been suggested that this capability of bacteria is dependent on a complex network of transcriptional interactions [9, 10]. These types of networks are called gene regulatory networks.
@@
The main feature of RC is that memory encoding and prediction are handled by two separate network substructures, as depicted in Fig. 1.
Figure 1: Structural and functional organization of the gene regulatory network of Escherichia coli. The genes depicted in green constitute a recurrent subgraph featuring cyclic paths (corresponding to the reservoir) capable of storing the recent past within its dynamics. The nodes in red form a directed acyclic subgraph (corresponding to the readout) responsible for interpreting the state of the reservoir. The network architecture shown here was obtained from the Ecocyc database [15]. The representation is adapted from Gabalda-Sagarra, Carey, and Garcia-Ojalvo [11].
@@
    It performs a nonlinear expansion of the incoming signals (input) in a space of high dimensionality.
@@
    It provides the system with memory by enabling it to retain information from its recent past.
The second substructure shown in Fig. 1 (red nodes) is the readout, located downstream of the reservoir and exhibiting a strictly feedforward architecture. This substructure combines the record of the immediate past encoded in the reservoir with the present input of the system, to generate predictions or perform classifications. This separation of roles allows to focus the training process (i.e. the modification of the network’s synaptic strengths so that the desired output is produced) solely on the readout [16].
In the classical RC approach, also known as an echo-state network (ESN), the reservoir is a random network with fixed (randomly set) weights, and solely the readout is trained. Since feedforward architectures are much easier to train than recurrent ones, the RC framework combines the computational power of a recurrent network with the training simplicity of a feedforward network [2]. RC has demonstrated outstanding performance across various benchmark tasks [17, 18, 19].
@@ II.1 Basic description
It is worth noting that the ESN approach requires tuning the network appropriately. Specifically, the recurrent subgraph should comply with the echo state property: its state $\mathbf{x}_{t}$ needs to be exclusively determined by the fading memory of the incoming signal $\mathbf{u}_{t}$ [17]. To put it differently, the reservoir state $\mathbf{x}_{t}$ should remain independent of the initial conditions that were present before the input $\mathbf{u}_{t}$, if we consider a long enough $\mathbf{u}_{t}$ signal. Empirically, it has been seen that if the spectral radius (defined as the maximum absolute eigenvalue of the reservoir connectivity matrix $W$) is smaller than unity, then the echo state property is satisfied. For large enough spectral radii, reservoirs exhibit multiple fixed points or periodic, even chaotic attractors (when the reservoir is sufficiently nonlinear). This violates the echo state property and prevent efficient training.
@@ II.2 Alternative training algorithms
To be operational, the learning algorithm must induce in the network a transition from chaotic to non-chaotic dynamics during training, and that is exactly what FORCE does. Specifically, Sussillo and Abbott [23] show that the chaotic dynamics during training can be quenched by using either external feedback loops (Fig. 3a) or internal synaptic modifications (Figs. 3b and 3c).
Figure 3: Network architectures for the FORCE algorithm [23]. In all cases, a reservoir (green nodes) drives a linear readout unit (red node). Weights are only changed for the connections highlighted in red. (a) The readout feeds back into the reservoir. (b) A separate recurrent network (smaller circled network) feeds back into the reservoir. (c) No additional feedback (beyond the recurrence in the reservoir) is needed, and all connections are trained.
All three schemes lead to the generation of intricate yet controlled outputs. Driven by any of these two strategies, the key for suppressing chaos is making strong and rapid synaptic modifications. In this regard, FORCE learning works in a manner distinct from traditional neural network training. Typically, conventional training involves a series of adjustments aimed at gradually diminishing initially large errors in the network output. The FORCE algorithm, in turn, reduces significantly the error’s size at the initial weight update, and maintains the errors small during the rest of the training procedure. In order to achieve this, the method makes use of the recursive least-squares (RLS) algorithm [24] as the update rule for the trained weights (see Sussillo and Abbott [23] for details). Further modification becomes unnecessary upon completion of the training, and the network can autonomously produce the desired output.
Finally, it is worth emphasizing that the FORCE algorithm allows us to train not only the readout weights of the network, such as in the original RC framework (Fig. 3a), but also RNNs with the architectures shown in Figs. 3b and 3c. In these architectures, training affects not only the readout weights but also the internal synaptic connections. These networks resemble are more realistic biologically than that of Fig. 3a, where feedback is only mediated by the readout unit. By enabling training of different architectures, the FORCE algorithm extends the original RC framework, and gives more flexibility and biological feasibility to the concept of computation through recurrence.
@@ II.3 Using dynamical systems theory to understand recurrence-based computations
Since supervised training dictates the computation to be performed without constraining how the network executes this computation, the exact method through which trained RNNs achieve their intended functions remains an unanswered question. In essence, the resulting networks have predominantly been regarded as black boxes. This differs from network models deliberately designed to implement specific established mechanisms, such as those proposed, for instance, by Hopfield [25] and Wang [26]. One way of “opening the black box”, proposed by Sussillo and Barak [27], is to treat the network as a nonlinear dynamical system.
Sussillo and Barak [27] explored not only the existence of stable and unstable fixed points and their linearized dynamics, but also the linearization in regions of phase space that are not genuine fixed points, but rather points where the system transitions very slowly (slow points). They also present a straightforward optimization algorithm for finding such areas. Linearization around each one of the fixed and slow points provides a mechanistic description of how high-dimensional RNNs implement their computations [27]. This final point is crucial for formulating hypotheses on how the task could potentially be solved by the biological network. Therein lies the importance of this dynamical-system based approach.
Optimization tools dependent on predefined analytic derivatives, such as those of the hyperbolic tangent function, are needed in order to find fixed and slow points of a RNN (see e.g. Section 3.2 of Sussillo and Barak [27] for details). Algorithms to find fixed points have been developed for simple (Elman) RNNs [28], and Hopfield-like RNN architectures [27, 29], but can be inconvenient for complicated models, or for studies that involve a combination of different RNN models. In this sense, a Tensorflow toolbox designed for determining fixed points and linearized dynamics in recurrent networks was introduced by Golub and Sussillo [30]. This software profits from TensorFlow’s highly efficient automatic differentiation back-end to execute the necessary optimizations and compute Jacobians, essential for the identification and characterization of fixed points.
@@ III.1 The mammalian cerebral cortex
Cortical circuits were inspirational to Maass, Natschläger, and Markram [14] when conceiving LSMs. It has been proposed that cortical networks are likely to operate following an RC paradigm, or at least that they incorporate most RC principles. One indication of this is the observable variation in both morphology and functionality among neurons across the cortex. Heterogeneity in neural characteristics has been observed in cortical neurons, the visual cortex and the retina, as well as in networks cultured in vitro [31, 32, 33, 34, 35, 36]. In this direction, RC-like computation can make use of a very heterogeneous dynamics, that would arise from this diversity of neural components. In addition, cortical networks present a highly recursive nature [37]. It is unclear, however, to what extent recurrence plays a role in computation by cortical circuits, and efforts are under way to address this issue [31, 32]. In what follows, we review current work in this direction.
@@
Dominey, Arbib, and Joseph [38] developed a RC-based model of the cortico-striatal system to shed light on the mechanisms underlying the behavior of the oculomotor system in different contexts [39]. The model consists of two sub-networks, following the established organization of the oculomotor system of primates in cortical and subcortical structures. The first sub-network corresponds to the prefrontal cortex, which is responsible for a variety of cognitive tasks ranging from decision making to the regulation of social behavior [40]. The second sub-network corresponds to the striatum, which is a subcortical area.
In the model of Dominey, Arbib, and Joseph [38], the prefrontal sub-network has fixed recurrent connections between its neurons; while synapses from this structure to the striatum are modifiable [41]. The neural activity in the prefrontal network generates unique patterns of activity in response to the sequential visual inputs. These state-encoding patterns are then linked to the appropriate outputs (representing the oculomotor movements) by the striatum.
Later, Dominey and Ramus [42] extended this model to the context of language processing in infants. Specifically, the authors showed that the model described previously (which is also known as a temporal recurrent network) simulates the capabilities of babies to learn the sequential and rhythmic structure of language. For instance, the network can effectively encode sequential and temporal patterns, which can in turn be linked to specific behaviors in the presence of appropriate stimuli. To reproduce the ability of babies to extract abstract structure from sequential inputs, the authors added a second recurrent network with short-term memory. These works illustrate how inherent representational abilities of the structure of language might emerge from a shared neural architecture, separate from what is needed for representing abstract structures.
Augmenting the cortico-striatal model discussed above with a neurophysiologically grounded language model enables the resulting system to learn grammatical structures [43, 44]. Furthermore, the efficacy of the learning algorithm can be increased by combining the model with an RC framework [45]. The detailed history and the developments of the cortico-striatal model was reviewed by Dominey [46].
@@
The visual system provides an excellent example of the need of processing time-dependent information. Nikolić et al. [33] obtained experimental evidence showing that the primary visual cortex of cats exhibits fading memory, which is one of the core features of ESNs. Specifically, in vivo spiking activities of 60–100 neurons were obtained for a sequence of different visual stimuli (letters). The authors demonstrated that the neuronal responses that they recorded at a given time had information about the letter presented at that moment but also, interestingly, about the letters presented right before. In other words, the population activity was shown to encode both the present and the recent past. As further verification, the authors used a network of linear integrate-and-fire neurons as a readout, to perform a basic linear classification that enabled them to identify various visual stimuli reliably from the spike trains of the recorded neurons (representing reservoir states) [33].
The emerging viewpoint discussed above suggests that the visual cortex can process sequential inputs by engaging in time-dependent computations that provide the cell with the ability to store its recent past. This stands in stark contrast to the traditional belief that the brain process information locally in time, on a frame-by-frame basis, through feedforward network architectures [47] that are not well suited to process dynamic visual scenes.
@@ III.2 Gene regulatory networks as reservoirs
Biological computation through recurrence does not necessarily need to arise in neural substrates, but can be supported by other biological networks. In this section, we review two example studies involving gene regulatory networks (GRNs). The first study looks into the ability of Escherichia coli to discriminate between time-dependent series of external chemical signals. The second one studies the memory capacities of the GRNs of five evolutionary distant organisms.
@@
Jones et al. [48] proposed that the gene regulatory network (GRN) of the bacterium Escherichia coli can act as a computational reservoir, receiving input signals from a sequence of chemical stimuli and acting upon a set of target proteins as an output. Specifically, the authors exposed a liquid culture of E. coli cells to sequences of inputs that represent different combinations of temperature and chemical conditions of the environment. Subsequently they studied the transcriptional state of the GRN using microarrays, to test whether the network was capable of making complex perceptual classification of the different inputs into desired classes. Numerical simulations of the aforementioned experimental setup showed that the RC framework could classify inputs with good performance, and perform logic operations such as XOR integration.
This work motivated an experimental study in which Didovyk et al. [49] employed genetic engineering techniques to manipulate bacterial cell cultures, aiming to create a range of basic classifiers capable of distinguishing between various chemical inputs. The outputs could be further combined to solve more intricate classification problems.
@@
The recurrence in the previous example was mostly caused by sparse self-regulation of a few genes, which was enough for the tasks proposed, which did not require memory. Fully recurrent topologies are necessary, however, for processing time-dependent information. In this sense, Gabalda-Sagarra, Carey, and Garcia-Ojalvo [11] studied the GRNs of five different organisms (from bacteria to humans) in standard memory-demanding tasks. Using public datasets, the authors identified that all five GRNs contained a recurrent reservoir formed by a small number of genes, while the majority of the genes in the networks were located downstream of the recurrent sub-graph, forming a feedforward network that can act as a readout, interpreting the high-dimensional dynamical state of the reservoir. In that way, the structure of GRNs are consistent with the echo-state network paradigm, providing the organism with a memory mechanism necessary to process temporal information.